A bank makes a 30 year Fully Amortizing FRM for $1,000,000 at an annual interest rate of 4.13% compounded monthly, with monthly payments. Suppose inflation is 2% per year, compounded monthly. What is the real value of the 20th payment?"
Answer #1

Monthly payment=Loan*(rate/12)/(1-1/(1+rate/12)^(12*loan term))=1000000*(4.13%/12)/(1-1/(1+4.13%/12)^(12*30))=4849.402724

Nominal 20th payment=4849.402724

Real 20th payment=4849.402724/(1+2%/12)^20=4690.550539